OpenStudy (anonymous):

Write an equation for the line in point slope form and then convert it to standard form. Slope: -3, Through (4,0)

OpenStudy (anonymous):

So far I have y-0=-3(x-4)

OpenStudy (skullpatrol):

Looks good so far :-)

OpenStudy (anonymous):

I know the next step is simplify, then write in standard form. I don't want to get it wrong though so Can you help with the next two steps?

OpenStudy (skullpatrol):

I can try.

OpenStudy (skullpatrol):

Recall the distributive property: a(b-c)=ab-ac

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

Skull im not even gonna lie lol Its 2:13 AM and Im kinda tired, Could you just show me the steps to getting the solution? but pick up from my part please.

OpenStudy (anonymous):

Any other time I would've wanted to go through everything but at this point I'd like the easy route lol.

OpenStudy (skullpatrol):

The "easy route" is just use what you already know about the distributive property.

OpenStudy (skullpatrol):

That is why you are taught to memorize this property :-)

OpenStudy (skullpatrol):



OpenStudy (anonymous):

After that other step I put, this I what I got next.. y=-3x-3+0

OpenStudy (skullpatrol):